A sample of helium gas occupies 12.4 L at 23°C and 0.956 atm. What volume will it occupy at 40°C and 0.956 atm?

A) 7.13 L

B) 13.1 L

C) 11.7 L

D) 21.6 L

E) none of these
Answer #1

According to Ideal gas equation

P1V1/T1 = P2V2/T2

P1 = 0.956 atm, V1 = 12.4 L, T1 = 23+273 = 300 K

P2 = 0.956 atm, V2 = ?, T2 = 40+273 = 313 K

(0.956*12.4/(300)) = (0.956*V2/(313))

V2 = 12.94 L

Almost equal to 13 L